(1 pt) Use the Laplace transform to solve the following initial value problem: y +-6y + 9y = 0 y0) = 2, y(0) = 1 First, us
Answer #1

Sol Given yl_ 6ylt9y=0 Y(0)=2, 7(0/=) dfylly = 3²y(s)-sy(0)-yo)) Lefylz= SY(S)-Y(0) {y} - 6 Lofyty + 9 Lofy}=0/0 -> s’y (s) –- A=2, -3A+B=-1) -3(2)+B=-1) B=-1)+6 TA=2, B=5 substitute in ean ③ we get Y(S) = + (-5) Taking inverse laplace transformation

Plz plz plz plz don't forget to Like